OpenStudy (anonymous):

Lois wrote the equation shown below. 4x + 12y = 4(x + 3y) Which property did Lois apply? commutative property of addition associative property of addition distributive property of addition addition property of equality

OpenStudy (anonymous):

I've never really been good with these.

OpenStudy (anonymous):

Its distributive property

OpenStudy (anonymous):

Alright. Thank you so much.

OpenStudy (nurali):

⇒a+(b+c)=(a+b)+c Associative ⇒a(b+c)=ab+ac Distributive ⇒a+b=b+a Commutative

OpenStudy (anonymous):

Thank you.

OpenStudy (nurali):

4x + 12y = 4(x + 3y) ab+ac=a(b+c) Distributive property of addition

OpenStudy (nurali):

My Pleasure.